Sandy pulls a boat a distance of 15.8 meters along a dock using a rope that makes an angle of 20.0o​ ​ with the horizontal. If she uses a force of 888 N, calculate work?
1
Expert's answer
2020-11-20T09:31:14-0500
"W=Fs\cos{20}\\W=(888)(15.8)\cos{20}=13200\ J"